OpenStudy (anonymous):

Simplify 5(4[3(2+x)])-5(-4[-3(2-x)])

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@gudden @aajugdar

OpenStudy (anonymous):

you just have to multiply and solve

OpenStudy (anonymous):

.. ._. I knew that.. but howwww

OpenStudy (anonymous):

@efco_chen

OpenStudy (anonymous):

@sammixboo

OpenStudy (gudden):

5(4[3(2+x)])-5(-4[-3(2-x)]) First of all multiply -4[-3(2-x)] Then multiply it with 5... Then subtract it from the following resultant 4[3(2+x)] multiplied by 5...

OpenStudy (gudden):

BTW.. did you notice something special in this question ?? It has got a super easy way of solving too..!

OpenStudy (anonymous):

12(-8-(-4x)) x 5?

OpenStudy (anonymous):

No, I just see a bunch of numbers clustered together with a bunch of barracks and parentheses.

OpenStudy (anonymous):

@mathmale

OpenStudy (gudden):

look at the question carefully 5(4[3(2+x)])-5(-4[-3(2-x)]) Can you see that first it has got something in common... ?

OpenStudy (gudden):

*first and second term have

OpenStudy (anonymous):

The 5.

OpenStudy (mathmale):

@kemist: You called? Exactly what do you need from me? The key to solving this problem lies in understanding "order of operations" rules. In this case, anything eiwthin parentheses has to be done first. Next, any multiplication has to be done BEFORE any addition / subtraction. Are you following these rules?

OpenStudy (gudden):

Alright 5 is correct.... Can you see that 4 and 3 are common too! And as mathmale said... Follow the BODMAS...

OpenStudy (anonymous):

PEMDAS Please excuse my dope a** swag. No I jk lol. Okay, let me try again.

OpenStudy (anonymous):

So are 2 and x @gudden

OpenStudy (gudden):

B = bracket O = open D = division M = Multiplication A = addition S = subtraction....

OpenStudy (anonymous):

^ I've never learned that

OpenStudy (anonymous):

P= Parentheses E= Exponents M= Multiply D= Divide A= Add S= Subtract

OpenStudy (gudden):

Well that's a single term..! ( ) any thing in this represents a single term..! And you cannot take that in common since they got different sings...! But if you notice the other numbers as I mentioned... i.e. 5,4 and 3.. are of opposite signs too! But we can leave a negative sign behind...^_^

OpenStudy (anonymous):

Okay, never knew that either. But I'm following.

OpenStudy (mathmale):

Order of operations rules are of critical importance. both in communicating methematical expressions to others and in interpreting mathematical expressions communicated to you, and you'll be on the same wavelength. It's only when these rules are not followed do we get into trouble, as in typing in mathematical expressions that could be interpreted in more than one way.

OpenStudy (anonymous):

5(4[3(2+x)])-5(-4[-3(2-x)]) 5(4[6+3x)])-5(-4[-6-x]) 5(24+12x)-5(-24+(-12x)) 120+60x-(-120)+(-60x)

OpenStudy (anonymous):

I don't know I'm just putting numbers together.

OpenStudy (anonymous):

iQuit

OpenStudy (gudden):

A little correction... 5(-4[-6-x]) Multiply the x with 3 too! And - 3*-x = +3x... ^_^ That's the only mistake.... Just correct it.. and your answer would be correct...^_^

OpenStudy (mathmale):

Focusing on part of the expression provided by Kemist: 5(4[3(2+x)]) Start on the inside and work your way out, of course following order of operations rules all the way. 3((2+x) must be done first. Why? Because it's inside brackets: [3(2+x)]. Result? 6+3x. Then we have 5(4(6+3x)). Inside the outer pair of parentheses we have 4(6+3x). This must be done now. Result: 24-12x. Now we have 5(24+12x). What next?

OpenStudy (anonymous):

So, what's the answer? ._. My head's starting to hurt.

OpenStudy (anonymous):

Move to the next brackets right?

OpenStudy (mathmale):

Now you try simplifying the 2nd term of the given expression. Sorry about your head, Kemist, but this is not a place to ask for answers. Please follow my example and try simplifying that second term. Or, take a break and come back to this later on.

OpenStudy (mathmale):

You try it. I'll comment. But I'd like to see you involved, doing something, even if you're not completely sure of what to do.

OpenStudy (mathmale):

We are now focusing on -5(-4[-3(2-x)]) only, and wish to simplify it.

OpenStudy (anonymous):

It's a test, can't take a break. Okay so 2nd term is -5(-4[-3(2-x)]) Like you said inside the brackets so [-3(2-x)] = -6-(-3x) = -5(-4(-6-3x) ?

OpenStudy (mathmale):

In -5(-4[-3(2-x)]), we have to focus on what's inside those square brackets first, yes. Try again on that. Watch your signs carefully.

OpenStudy (anonymous):

That's what I was doing. My numbers are wrong or the signs? @mathmale

OpenStudy (anonymous):

@iGreen

OpenStudy (gudden):

The signs are going wrong...

OpenStudy (anonymous):

Okay -5(-4(-6+3x)?

OpenStudy (mathmale):

Start over. Focus on the innermost brackets and do the work INSIDE those brackets first, making sure that your resulting signs are correct.

OpenStudy (anonymous):

-5(-4[-3(2-x)]) = [-3(2-x)] = -6+3x = -5(-4(-6+3x) ?

OpenStudy (anonymous):

5(24+12x)-5(-4(-6+3x) = 5(24+12x)-5(24+(-12x)

OpenStudy (mathmale):

Better. You had a sign error before. to keep matters straight, let's do this problem step by step. Focus on the innermost brackets / parentheses: -3(2-x). Nothing can be done inside those parentheses, so move on to what's inside the square brackets: -3(2-x)]. Simplify that, please.

OpenStudy (anonymous):

120+60x-120+60x FINAL ANSWER D:

OpenStudy (anonymous):

Okay, -6-3x

OpenStudy (mathmale):

No. Sign error, again. Try again, please.

OpenStudy (anonymous):

-6+3x

OpenStudy (mathmale):

Better. See why I wanted you to do this problem step by step? Then -5(-4[-3(2-x)]) becomes -5(-4[-6+3x]). What next? Explain in words what's next, before you actually do the work.

OpenStudy (anonymous):

Multiply the -4 with the numbers in parentheses?

OpenStudy (anonymous):

Right? @mathmale

OpenStudy (mathmale):

Yes. I'd phrase it this way: "Multiply the quantity within the square brackets by -4." do that now, please. Hold the -5( ) until later.

OpenStudy (anonymous):

-5(24-12x)

OpenStudy (anonymous):

Right or wrong?

OpenStudy (anonymous):

@SkylarLee Do this.